OpenStudy (anonymous):

please help!! medals given rewrite the expresion cos^4x as an equivalent expression that does not contain powers of trigonometric functions greater than 1

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@asnaseer

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@phi

OpenStudy (phi):

\[ \cos^2(x) = \frac{1}{2}(1 + \cos(2x)) \] \[ \cos^4x = \cos^2 x \cdot \cos^2 x = \frac{1}{2}(1 + \cos(2x))\frac{1}{2}(1 + \cos(2x)) \]

OpenStudy (phi):

you can multiply out \[ \frac{1}{2}(1 + \cos(2x))\frac{1}{2}(1 + \cos(2x)) = \frac{1}{4} (1+\cos 2x)(1+\cos 2x) \] and then try simplifying again

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

\[\frac{ 1 }{ 4}(1+\cos(4x)\]

OpenStudy (anonymous):

what do i do next?

OpenStudy (anonymous):

@Directrix